OpenStudy (anonymous):

Can someone help me with this. Find the limit of this problem limit as x-->0 (³√(1+2x)-1)/(x)

OpenStudy (anonymous):

is that a cube root inclusive of -1 term in numearator, or is -1 out of numerator

OpenStudy (anonymous):

its out of the numerator

OpenStudy (anonymous):

i guess what i am trying to say, is does that cube root include the -1 term

OpenStudy (anonymous):

³√(1+2x)-1

OpenStudy (anonymous):

no it should be separate. sorry ³√(1+2x) - 1

OpenStudy (anonymous):

limit does not exist

OpenStudy (anonymous):

my textbook says it was 2/3.

OpenStudy (anonymous):

well then can you please retype the question,

OpenStudy (anonymous):

heres what I'd suggest multiply numerator and denominator by conjugate

OpenStudy (anonymous):

(³√(1+2x)-1)/ x

OpenStudy (anonymous):

multiply by that cube root term + 1 and in numerator you will have to differnce of two squares concept

OpenStudy (anonymous):

which will give you 1+2x-1 and in denominator you will have (x)* ( cuberoot term +1)

OpenStudy (anonymous):

no cancel out that x with the x in denominator

OpenStudy (anonymous):

finally put x=0 , i.e evaluate limits as you cant simply further

OpenStudy (anonymous):

so i got this: (1+2x-1)/ (x (³√(1+2x)+1))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cancel x from num and den

OpenStudy (anonymous):

you said cancel out the x so, is it like this? (1+2-1)/ (³√(1+2x)+1)

OpenStudy (anonymous):

then, 2/ (³√(1+2x)+1)

OpenStudy (anonymous):

but if i plug in 0, it gives me 1. i don't know what i did wrong.

OpenStudy (anonymous):

okie, i guess you must retype the entire question

OpenStudy (anonymous):

there seems to be a minor error in typing those brackets

OpenStudy (anonymous):

(³√(1+2x)- 1)/ x

OpenStudy (anonymous):

one moment, are you allowed to use L' hospital rule ?

OpenStudy (anonymous):

cube root doesn't involve minus1

OpenStudy (anonymous):

i haven't learned that yet

OpenStudy (anonymous):

the correct factor to simplify this with is \[\frac{ \sqrt[2/3]{1+2x} +1 }{\sqrt[2/3]{1+2x} +1}\]

OpenStudy (anonymous):

because a simple look and it tells you that this qustion is of 0/0 form meaning you need to use it

OpenStudy (anonymous):

that will give you 2/3

OpenStudy (anonymous):

could u tell me what L-Hopsital rule is?

OpenStudy (anonymous):

and Algebraic, how did you get that?

OpenStudy (anonymous):

algebraic is incorrect, you cant use perfect square

OpenStudy (anonymous):

as it 0/0 form

OpenStudy (anonymous):

put x =0 and you will find numerator changes to 1+0-1 and denominator -=0

OpenStudy (anonymous):

ok i get that. but how is it 2/3 then?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

hey algebraic do you know how to do this?

OpenStudy (anonymous):

please. i need this

OpenStudy (anonymous):

anyone still out there?

OpenStudy (anonymous):

fml

OpenStudy (anonymous):

5 more mintues...

OpenStudy (anonymous):

if any1 is listening to this, i've only got a few minutes left.

OpenStudy (anonymous):

damn i got to go... last chance any1?

OpenStudy (anonymous):

Guess not

OpenStudy (anonymous):

psi or algebraic, any of you still there?

OpenStudy (anonymous):

last chace guys!!!

OpenStudy (anonymous):

thats it screw this place.

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

whelp, the way I mentioned works, but it's long. I stumbled across a simple way while playing with it however.

OpenStudy (anonymous):

make a substitution u=2x+1 then x=(u-1)/2 and your expression looks like \[2*\frac{ u ^{1/3} -1 }{u-1}\]

OpenStudy (anonymous):

the denom. is a difference of cubes : u^(1/3) and 1

OpenStudy (anonymous):

use the difference of cubes identity to expand that

OpenStudy (anonymous):

and the numerator becomes 2 and after taking the limit the denom becomes 3:)

OpenStudy (anonymous):

Apply L'Hospital's rule. Using Mathematica notation below. D[f(x),x] means take the derivative of f(x) with respect to x. \[\text{Limit}\left[\frac{D\left[\sqrt[3]{1+2x}-1,x\right]}{D[x,x]},x\to 0\right]\to \text{Limit}\left[\frac{\frac{2}{3 (1+2 x)^{2/3}}}{1},x\to 0\right] \]

OpenStudy (anonymous):

don't think he's allowed to. thanks for bringing up l'Hopital's for the hundredth time though.

OpenStudy (anonymous):

well thanks for your help.

OpenStudy (anonymous):

did you follow my explanation @abc14 ?

OpenStudy (anonymous):

yeah i think i know how to do it now

OpenStudy (anonymous):

a^3 – b^3 = (a – b)(a^2 + ab + b^2) a= u^(1/3) b=1

OpenStudy (anonymous):

try it :)

OpenStudy (anonymous):

((U^1/3)-1)(u^2/3 + u^1/3 +1) is tthat correct algebraic

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

that looks ugly tho

OpenStudy (anonymous):

naw

OpenStudy (anonymous):

numerator was 2*(u^(1/3) -1)

OpenStudy (anonymous):

so what cancels?

OpenStudy (anonymous):

2((U^1/3) -1)(u^2/3 + u^1/3 +1) i don't see anything that can be

OpenStudy (anonymous):

(u^(1/3) -1)

OpenStudy (anonymous):

2*(u^(1/3) -1) ______________________________________ (u^1/3)-1)(u^2/3 + u^1/3 +1)

OpenStudy (anonymous):

so 2 over (u^2/3 + u^1/3 +1)

OpenStudy (anonymous):

lim (u^2/3 + u^1/3 +1) = 3 lol as x->0

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

u^2/3 + u^1/3 = u?

OpenStudy (anonymous):

\[2* \frac{ u ^{1/3} -1 }{ (u ^{1/3} -1)(u ^{2/3} +u ^{1/3} +1)}\]

OpenStudy (anonymous):

\[\frac{ 2 }{u ^{2/3} +u ^{1/3} +1}\]

OpenStudy (anonymous):

put x's back in for your u's \[\frac{ 2 }{( (2x+1)^{2/3} +(2x+1)^{1/3} +1}\]

OpenStudy (anonymous):

take the limit as x->0

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

thank you algebraic!

OpenStudy (anonymous):

no problem:)